OpenStudy (anonymous):

simplify 3 to the power of 3 square root 375

OpenStudy (anonymous):



OpenStudy (anonymous):

do you have an idea on how to do this? if you try it I can help you with any problems

OpenStudy (anonymous):

i have no idea how to solve this

OpenStudy (anonymous):

okay well we need to to find factors of 375, what are they?

OpenStudy (anonymous):

Any idea how surds work at all?

OpenStudy (anonymous):

no..

OpenStudy (anonymous):

375=3*5^3

OpenStudy (anonymous):

that's okay when we simplify we want to separate the square root into this \[\sqrt{x^2}=x\] For example is I have \[\sqrt{75}=\sqrt{3*25}\]\[=\sqrt{3}\sqrt{25}\]\[=\sqrt{3}\sqrt{5^2}\]\[=5 \sqrt{3}\]

OpenStudy (anonymous):

cherrypot try this one \[\sqrt{72}\]

OpenStudy (anonymous):

@will why it is 135sqrt(6)

OpenStudy (anonymous):

should be \[135\sqrt{15}\]

OpenStudy (anonymous):

\[3^2\sqrt{375}=9\sqrt{25*15}\]\[=9\sqrt{25}\sqrt{15}\]\[=9*5\sqrt{15}\]\[=45\sqrt{15}\]

OpenStudy (anonymous):

how are you going with the \[\sqrt{72}\] cherrypot?

OpenStudy (anonymous):

It's 3 cubed Zed, need to multiply by 3 again

OpenStudy (anonymous):

my apologies i thought it said squared. it is \[135\sqrt{15}\] thanks Will!

OpenStudy (anonymous):

i think it is cubic root?

OpenStudy (anonymous):

3 power 3, sq. root 375

OpenStudy (anonymous):

it says square root in the original post, Yifan :)

OpenStudy (anonymous):

oh..i ididnt see that